OpenStudy (anonymous):

.Solve: log (2x + 1) = log (x - 1).

OpenStudy (unklerhaukus):

exponentiate both sides

OpenStudy (unklerhaukus):

\[\begin{align}\log (2x + 1) &= \log (x - 1)\\e^{\log (2x + 1) }&= e^{\log (x - 1)}\\2x+1&=x-1\end{align}\]

OpenStudy (anonymous):

then you solve for x?

OpenStudy (unklerhaukus):

yeah, what do you get?

OpenStudy (anonymous):

x=-2

OpenStudy (unklerhaukus):

if you would like to check your solution \[\begin{align}\log (2(-2) + 1) &= \log ((-2) - 1)\\\log(-3)&=\log(-3)\end{align}\] which is always true

OpenStudy (unklerhaukus):

so you have found the solution for x